OpenStudy (anonymous):

how many solutions does 5x+48+7x=12(x+4)

OpenStudy (anonymous):

WHen you see x to the first power, without any additional powers, it is usually 1 solution.

OpenStudy (anonymous):

in this case it is infinite solutions, though.

OpenStudy (anonymous):

5x+48+7x=12(x+4) 5x+48+7x=12x+48 12x+48=12x+48

OpenStudy (anonymous):

in this case it is infinite solutions, though.

OpenStudy (anonymous):

WHen you see x to the first power, without any additional powers, it is usually 1 solution.